OpenStudy (anonymous):

solve the following equation for x: logx + log(x-3) = log(3x) A. x = 6 B. x = 0,6 C. x = 0 D. There are no solutions.

jimthompson5910 (jim_thompson5910):

log(x) + log(x-3) = log(3x) log(x(x-3)) = log(3x) x(x-3) = 3x x^2-3x = 3x x^2-3x-3x = 0 x^2-6x = 0 I'll let you finish. Don't forget to check your solutions.